OpenStudy (anonymous):

Solve by substitution method 2x-5y=-1 x -3y=-3 I really want step by step please.

OpenStudy (anonymous):

2x-5y=-1 x-3y=-3 x-3y=-3 we solve for x. Since it's easier. we do this by adding -3y to both sides X=-3+3y then we plug in (-3+3y) in for x in the equation 2x-5y+=-1 2(-3+3y)-5y=-1 We distributed the 2(-3+3y) it equals -6y+6y in which the cancel out each other. we get -5y=-1 Divide both sides by -5 \[y=\frac{1}{5}\]

OpenStudy (anonymous):

thanks I still dont understand

OpenStudy (anonymous):

oops sorry i made a mistake. i'll correct it right now.

OpenStudy (anonymous):

2x-5y=-1 x-3y=-3 x-3y=-3 we solve for x. Since it's easier. we do this by adding -3y to both sides X=-3+3y then we plug in (-3+3y) in for x in the equation 2x-5y+=-1 2(-3+3y)-5y=-1 We distributed the 2(-3+3y) it equals -6+6y -6+6y-5y=-1 6y-6y =y -6+y=-1 subtract negative -6 from both sides. -1--6=5 y=5